# MoH: Multi-Head Attention as Mixture-of-Head Attention

### v1

@@ Abstract
In this work, we upgrade the multi-head attention mechanism, the core of the Transformer model, to improve efficiency while maintaining or surpassing the previous accuracy level. We show that multi-head attention can be expressed in the summation form. Drawing on the insight that not all attention heads hold equal significance, we propose Mixture-of-Head attention (MoH), a new architecture that treats attention heads as experts in the Mixture-of-Experts (MoE) mechanism. MoH has two significant advantages: First, MoH enables each token to select the appropriate attention heads, enhancing inference efficiency without compromising accuracy or increasing the number of parameters. Second, MoH replaces the standard summation in multi-head attention with a weighted summation, introducing flexibility to the attention mechanism and unlocking extra performance potential. Extensive experiments on ViT, DiT, and LLMs demonstrate that MoH outperforms multi-head attention by using only 50$\sim$90of the attention heads. Moreover, we demonstrate that pre-trained multi-head attention models, such as LLaMA3-8B, can be further continue-tuned into our MoH models. Notably, MoH-LLaMA3-8B achieves an average accuracy of 64.0% across 14 benchmarks, outperforming LLaMA3-8B by 2.4% by utilizing only 75% of the attention heads. We believe the proposed MoH is a promising alternative to multi-head attention and provides a strong foundation for developing advanced and efficient attention-based models.
@@ 1 Introduction
Since attention is introduced and becomes a fundamental component of Transformers (Vaswani et al., 2017), multi-head attention has been the standard architecture for natural language processing (Kenton & Toutanova, 2019) and computer vision tasks (Dosovitskiy et al., 2021). It is well known that using multiple heads can improve model accuracy. However, not all attention heads hold equal significance. Some works have shown that many attention heads can be pruned without affecting accuracy. For example, Voita et al. (2019) introduces a method to quantify the usefulness of each attention head and prune those that are redundant. Similarly, Michel et al. (2019) challenges the necessity of multiple heads by examining the impact of extensive pruning across various settings. These findings demonstrate that vanilla multi-head attention contains redundant attention heads.
@@
Besides, in multi-head attention, each attention head operates in parallel, and the final output is the sum of all attention heads (please refer to Section 3.1). Given that these attention heads operate independently and some may be redundant, we argue that it is possible to build a dynamic attention-head routing mechanism. Such a mechanism would enable each token to adaptively select the appropriate attention heads, enhancing inference efficiency without compromising accuracy.
@@
To this end, we introduce Mixture-of-Head attention (MoH), a new architecture that integrates multi-head attention with the Mixture-of-Experts (MoE) mechanism (Jacobs et al., 1991; Jin et al., 2024b). Specifically, we propose to treat attention heads as experts within the MoE framework. Similar to MoE, MoH consists of multiple attention heads and a router that activates the Top-K heads for each token. Moreover, we replace the standard summation in multi-head attention with a weighted summation. This design offers two significant advantages: First, MoH allows each token to select the most relevant attention heads, improving inference efficiency without sacrificing accuracy or increasing the parameters. Second, by replacing the standard summation in multi-head attention with a weighted summation, MoH enhances the flexibility of the attention mechanism and increases the performance potential. Moreover, to efficiently capture common knowledge across different contexts, we designate a subset of attention heads as shared heads that remain always activated.
@@
We evaluate our proposed MoH across various popular model frameworks, including Vision Transformers (ViT) (Dosovitskiy et al., 2021) for image classification, Diffusion models with Transformers (DiT) (Peebles & Xie, 2023) for class-conditional image generation, and Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022) for language tasks. We show that MoH achieves competitive performance, or even outperforms multi-head attention with only 50$\sim$90of the attention heads. For example, MoH-ViT-B achieves 84.9%/84.7% Top-1 accuracy on the ImageNet-1K (Deng et al., 2009) classification benchmark, surpassing well-tuned multi-head attention baselines with only 75%/50% of the attention heads.
@@
  • A wide range of experiments across various popular model frameworks, including ViT, DiT, and LLMs, confirm that MoH is a promising alternative to vanilla multi-head attention, laying a solid foundation for developing advanced and efficient attention-based models.
@@ 2 Related Work
Mixture-of-Experts Models.    The Mixture-of-Experts (MoE) method (Du et al., 2022; Lewis et al., 2021; Rajbhandari et al., 2022; Roller et al., 2021; Zhou et al., 2022; Jin et al., 2024b) is introduced to expand the capacity of deep neural networks without increasing computational costs. In this approach, only a subset of parameters, known as experts, is activated for each input. Shazeer et al. (2017) first introduces an MoE layer between LSTM layers. Switch Transformer (Fedus et al., 2022) further simplifies the gating mechanism by selecting only the Top-1 expert per token. Gshard (Lepikhin et al., 2021) improves the Top-2 expert routing strategy. In contrast to MoE, which emphasizes efficient parameter scaling while maintaining manageable computational costs, the proposed MoH focuses on reducing the activation of redundant attention heads without increasing the number of parameters.
Figure 1: A high-level comparison between the multi-head attention and our proposed mixture-of-head attention. Subfigure (a) illustrates a standard multi-head attention layer with $h$ attention heads, while subfigure (b) demonstrates the Mixture-of-Head attention (MoH) architecture. It is important to note that MoH does not increase the number of attention heads, ensuring that the total parameter for MoH is comparable to that of the multi-head attention.
@@ 3 Methodology
In this work, we aim to reduce the activation of redundant attention heads without increasing the number of parameters. A high-level comparison between the vanilla multi-head attention and our proposed Mixture-of-Head attention (MoH) is presented in Fig. 1.
@@ 3.1 Multi-Head Attention
We begin by reviewing the standard multi-head attention mechanism introduced by Vaswani et al. (2017). The multi-head attention mechanism is based on scaled dot-product attention. Specifically, for $T$ tokens $\bm{X}\in\mathbb{R}^{T\times d_{in}}$ of $d_{in}$ dimensions each and $T^{\prime}$ tokens $\bm{X}^{\prime}\in\mathbb{R}^{T^{\prime}\times d_{in}}$ of $d_{in}$ dimensions each, the scaled dot-product attention is computed as follows:
@@
Concatenation Form.    To enhance the representation power of the attention layer, Vaswani et al. (2017) proposes to allow multiple attention heads to operate on different low-dimensional projections of the input tokens. Specifically, the multi-head attention mechanism computes $h$ different low-dimensional projections of $(\bm{Q},\bm{K},\bm{V})$, performs scaled dot-product attention for each head, concatenates the results, and applies a final projection to the concatenated output. The concatenation form of the multi-head attention can be formulated as:
@@
where $\bm{W}_{Q}^{i}\in\mathbb{R}^{d_{in}\times d_{k}/h}$, $\bm{W}_{K}^{i}\in\mathbb{R}^{d_{in}\times d_{k}/h}$, and $\bm{W}_{V}^{i}\in\mathbb{R}^{d_{in}\times d_{v}/h}$ represent the $i_{th}$ projection matrices for the query, key, and value, respectively. $\bm{W}_{O}\in\mathbb{R}^{d_{v}\times d_{out}}$ is the final projection matrix.
@@
3.2 Mixture-of-Head Attention
Recently, the Mixture-of-Experts (MoE) method has emerged as a popular approach for scaling the parameters of large language models (Jiang et al., 2024). A typical MoE layer consists of multiple expert networks and a router that activates the Top-K experts. Generally, the number of activated experts $K$ is significantly smaller than the total number of experts to ensure inference efficiency.
Heads as Experts.    Inspired by the great success of MoE, we propose Mixture-of-Head attention (MoH), which treats attention heads as experts. Specifically, MoH consists of $h$ heads $\bm{H}=\{H^{1},H^{2},...,H^{h}\}$ and a router that activates the Top-K heads. Formally, given input tokens $\bm{X}$ and $\bm{X}^{\prime}$, the output of MoH is the weighted sum of outputs from the $K$ selected heads:
where $g_{i}$ represents the routing score. $g_{i}$ is non-zero only when the $i_{th}$ attention head is activated. This design provides two key advantages: On the one hand, MoH enables each token to select the most relevant attention heads, boosting inference efficiency while maintaining accuracy. On the other hand, in contrast to the standard summation in multi-head attention, the weighted summation in MoH enhances the flexibility of the attention mechanism and unlocks performance potential.
$\quad g_{i}=\begin{cases}\alpha_{1}\text{Softmax}(\bm{W}_{s}\bm{x}_{t})_{i},&\text{if}\ 1\leq i\leq h_{s},\\ \alpha_{2}\text{Softmax}(\bm{W}_{r}\bm{x}_{t})_{i},&\text{if}\ (\bm{W}_{r}\bm{x}_{t})_{i}\in\text{Top-K}\big{(}\{(\bm{W}_{r}\bm{x}_{t})_{i}|h_{s}+1\leq i\leq h\}\big{)},\\ 0,&\text{otherwise},\end{cases}$ (5)
where $h_{s}$ denotes the number of shared heads. $\bm{W}_{s}\in\mathbb{R}^{h_{s}\times d_{in}}$ and $\bm{W}_{r}\in\mathbb{R}^{(h-h_{s})\times d_{in}}$ represent the projection matrices for the shared and routed heads, respectively. The coefficients $\alpha_{1}$ and $\alpha_{2}$ balance the contributions of the shared and routed heads, and are defined as:
$[\alpha_{1},\alpha_{2}]=\text{Softmax}(\bm{W}_{h}\bm{x}_{t}),$ (6)
$\mathcal{L}_{b}=\sum_{i=h_{s}+1}^{h}f_{i}P_{i},\ f_{i}=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}(\text{Token $\bm{x}_{t}$ selects Head $i$}),\ P_{i}=\frac{1}{T}\sum_{t=1}^{T}\text{Softmax}(\bm{W}_{r}\bm{x}_{t})_{i},$ (7)
$\mathcal{L}=\mathcal{L}_{task}+\beta\mathcal{L}_{b},$ (8)
Table 1: Comparisons to current state-of-the-art methods on ImageNet-1K classification. All models are trained exclusively on the ImageNet-1K training set. Our MoH-ViT models, based on TransNeXt (Shi, 2024), are trained for 300 epochs using a resolution of 224$\times$224. To ensure a fair comparison, we only replace the standard multi-head attention with our Mixture-of-Head attention (MoH), keeping all other training parameters identical to TransNeXt.
@@ 3.2 Mixture-of-Head Attention
4 Experiments
4.1 ViT for Image Classification
Results.    As shown in Tab. 1, despite activating only a subset of attention heads, MoH-ViT achieves highly competitive performance compared to current state-of-the-art methods. For example, MoH-ViT-B achieves 84.9% Top-1 accuracy on the ImageNet-1K classification benchmark with just 75% of the attention head. In contrast, the well-established ViT baseline, TransNeXt, attains a slightly lower accuracy of 84.8% while requiring 100% of the heads to be activated. Tab. 1 demonstrates that MoH-ViT outperforms other models with fewer activated attention heads. This suggests that MoH is a promising alternative to vanilla multi-head attention for vision model design, offering the potential for competitive performance with more efficient attention head usage.
Table 2: Comparisons to DiT on the benchmarking of class-conditional image generation on ImageNet-1K at 256$\times$256 resolution. To ensure a fair comparison, we only replace the standard multi-head attention with the MoH in MoH-DiT models, while keeping all other training parameters identical to DiT. “400K” denotes the training budget is 400K training steps.
@@ 4.1 ViT for Image Classification
Table 3: Comparisons to current state-of-the-art methods on the benchmarking of class-conditional image generation on ImageNet-1K at 256$\times$256 resolution.$\uparrow$” denotes that higher is better. $\downarrow$” denotes that lower is better. “cfg” denotes the classifier-free diffusion guidance scale. We extend the training budget of our MoH-DiT-XL/2 to 7, [^0] K training steps, aligning it with DiT-XL/2.
[…]
 Heads (%)
ADM-G, ADM-U (Dhariwal & Nichol, 2021) - 3.94 6.14 215.84 0.83 0.53
CDM (Ho et al., 2022) - 4.88 - 158.71 - -
LDM-8 (Rombach et al., 2022) - 15.51 - 79.03 0.65 0.63
LDM-4 (Rombach et al., 2022) - 10.56 - 103.49 0.71 0.62
LDM-4-G (cfg=1.25) - 3.95 - 178.22 0.81 0.55
DiT-XL/2 7, [^0] K (Peebles & Xie, 2023) 100 9.62 6.85 121.50 0.67 0.67
DiT-XL/2 7, [^0] K (cfg=1.25) 100 3.22 5.28 201.77 0.76 0.62
MoH-DiT-XL/2 2, [^0] K 75 10.95 6.19 106.69 0.67 0.66
MoH-DiT-XL/2 2, [^0] K 90 10.67 6.15 107.80 0.67 0.65
MoH-DiT-XL/2 7, [^0] K 90 8.56 6.61 129.54 0.68 0.67
MoH-DiT-XL/2 7, [^0] K (cfg=1.25) 90 2.94 5.17 207.25 0.77 0.63
4.2 DiT for Class-Conditional Image Generation
Model Settings.    For Diffusion models with Transformers (DiT) (Peebles & Xie, 2023), we only replace the standard multi-head attention with our MoH in MoH-DiT models, while keeping all other training parameters identical to DiT. We use the ImageNet-1K dataset (Deng et al., 2009) for class-conditional image generation at a resolution of 256$\times$256.
Evaluation Benchmarks.    To evaluate generation performance, we use Frechet Inception Distance (FID(Heusel et al., 2017) to assess overall sample quality, Precision and Recall (Kynkäänniemi et al., 2019) to measure fidelity and diversity separately, and sFID (Nash et al., 2021) as a metric that better captures spatial relationships than FID. Moreover, we use Inception Score (IS(Salimans et al., 2016) as another metric for fidelity.
Results.    To conduct comparative evaluations of our proposed MoH-DiT models against vanilla DiT models, we start with Small models and expand to XLarge models. As shown in Tab. 2, MoH-DiT models consistently outperform vanilla DiT models with 90% of attention heads activated. However, when only 75% of the attention heads are activated, MoH-DiT models perform worse than DiT models with 100% of attention heads activated. This may be because image generation tasks are dense prediction tasks that require attention mechanisms to capture pixel-level fine-grained relationships, leaving less redundancy in the attention heads compared to image classification tasks. Moreover, we extend the training budget of our MoH-DiT-XL/2 to 7, [^0] K training steps, aligning it with DiT-XL/2. As shown in Tab. 3, despite activating 90% attention heads, MoH-DiT-XL/2 achieves highly competitive performance compared to current state-of-the-art methods. These results suggest that MoH is a promising alternative to multi-head attention for diffusion models.

### v2

@@ Abstract
In this work, we upgrade the multi-head attention mechanism, the core of the Transformer model, to reduce computational costs while maintaining or surpassing the previous accuracy level. We show that multi-head attention can be expressed in the summation form. Drawing on the insight that not all attention heads hold equal significance, we propose Mixture-of-Head attention (MoH), a new architecture that treats attention heads as experts in the Mixture-of-Experts (MoE) mechanism. MoH has two significant advantages: First, MoH enables each token to select the appropriate attention heads, enhancing inference efficiency without compromising accuracy or increasing the number of parameters. Second, MoH replaces the standard summation in multi-head attention with a weighted summation, introducing flexibility to the attention mechanism and unlocking extra performance potential. Extensive experiments on ViT, DiT, and LLMs demonstrate that MoH outperforms multi-head attention by using only 50$\sim$90of the attention heads. Moreover, we demonstrate that pre-trained multi-head attention models, such as LLaMA3-8B, can be further continue-tuned into our MoH models. Notably, MoH-LLaMA3-8B achieves an average accuracy of 64.0% across 14 benchmarks, outperforming LLaMA3-8B by 2.4% by utilizing only 75% of the attention heads. We believe the proposed MoH is a promising alternative to multi-head attention and provides a strong foundation for developing advanced and efficient attention-based models. The code is available at https://github.com/SkyworkAI/MoH.
Machine Learning, ICML
@@ 1 Introduction
Since attention is introduced and becomes a fundamental component of Transformers (Vaswani et al., 2017), multi-head attention has been the standard architecture for natural language processing (Kenton & Toutanova, 2019) and computer vision tasks (Dosovitskiy et al., 2021). It is well known that using multiple heads can improve model accuracy. However, not all attention heads hold equal significance. Some works have shown that many attention heads can be pruned without affecting accuracy. For example, Voita et al. (2019) introduces a method to quantify the usefulness of each attention head and prune those that are redundant. Similarly, Michel et al. (2019) challenges the necessity of multiple heads by examining the impact of extensive pruning across various settings. In computer vision, some works also identify attention head redundancy. Bhattacharyya et al. (2023) reduces redundancy to boost performance, while Yun & Ro (2024) develop single-head attention for efficiency. These findings demonstrate that vanilla multi-head attention contains redundant attention heads.
Figure 1: A high-level comparison between the multi-head attention and our proposed mixture-of-head attention. Subfigure (a) illustrates a standard multi-head attention layer with $h$ attention heads, while subfigure (b) demonstrates our proposed Mixture-of-Head attention (MoH) architecture. It is important to note that MoH does not increase the number of attention heads, ensuring that the total parameter for MoH is comparable to that of the multi-head attention.
@@
Besides, in multi-head attention, each head operates in parallel, and the final output is the sum of all heads (please refer to Section 3.1). Given that these attention heads operate independently and some may be redundant, we argue that it is possible to build a dynamic attention-head routing mechanism. Such a mechanism would enable each token to adaptively select the appropriate attention heads, enhancing inference efficiency without compromising accuracy.
@@
To this end, we introduce Mixture-of-Head attention (MoH), a new architecture that integrates multi-head attention with the Mixture-of-Experts (MoE) mechanism (Jacobs et al., 1991). Specifically, we propose to treat attention heads as experts within the MoE framework. Similar to MoE, MoH consists of multiple attention heads and a router that activates the Top-K heads for each token. Moreover, we replace the standard summation in multi-head attention with a weighted summation. This design offers two significant advantages: First, MoH allows each token to select the most relevant attention heads, improving inference efficiency without sacrificing accuracy or increasing the parameters. Second, by replacing the standard summation in multi-head attention with a weighted summation, MoH enhances the flexibility of the attention mechanism and increases the performance potential. Moreover, to efficiently capture common knowledge across different contexts, we designate a subset of attention heads as shared heads that remain always activated.
@@
We evaluate our proposed MoH across various popular model frameworks, including Vision Transformers (ViT) (Dosovitskiy et al., 2021) for image classification, Diffusion models with Transformers (DiT) (Peebles & Xie, 2023) for class-conditional image generation, and Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022). We show that MoH achieves competitive performance, or even outperforms multi-head attention with only 50$\sim$90of the attention heads. For example, MoH-ViT-B achieves 84.9%/84.7% Top-1 accuracy on the ImageNet-1K (Deng et al., 2009) classification benchmark, surpassing well-tuned multi-head attention baselines with only 75%/50% of the attention heads.
@@
  • Extensive experiments across various popular model frameworks, including ViT, DiT, and LLMs, confirm that MoH is a promising alternative to vanilla multi-head attention, laying a solid foundation for developing advanced and efficient attention-based models.
@@ 2 Related Work
Mixture-of-Experts Models.    The Mixture-of-Experts (MoE) method (Du et al., 2022; Lewis et al., 2021; Rajbhandari et al., 2022; Roller et al., 2021; Zhou et al., 2022; Jin et al., 2025) is introduced to expand the capacity of deep neural networks without increasing computational costs. In this approach, only a subset of parameters, known as experts, is activated for each input. Shazeer et al. (2017) first introduces an MoE layer between LSTM layers. Switch Transformer (Fedus et al., 2022) further simplifies the gating mechanism by selecting only the Top-1 expert per token. Gshard (Lepikhin et al., 2021) improves the Top-2 expert routing strategy. In contrast to MoE, which emphasizes efficient parameter scaling while maintaining manageable computational costs, our MoH focuses on reducing the activation of redundant attention heads without increasing the number of parameters.
Attention Head Specialization and Efficiency.    Many recent studies show that not all attention heads in Transformers are equally useful. Peng et al. (2020) proposes a mixture-of-heads approach, where only a few selected heads are used, yet the model performs just as well or even better. Csordás et al. (2024) pushes this further with SwitchHead, an MoE-style method that activates only a small number of heads for each token, speeding up inference while keeping performance high. In long-context language models, this idea is even more clear. Wu et al. (2024) shows that a few special retrieval heads are mainly responsible for keeping facts consistent in long inputs. Fu et al. (2024) finds that keeping only the most useful heads in the KV cache can save memory. Xiao et al. (2024) proposes DuoAttention, which combines different types of heads to make long-context inference more efficient without losing quality. Similar patterns appear in vision models. Gandelsman et al. (2023) shows that CLIP’s attention heads each focus on specific visual features, and this can be explained through related text prompts. Balasubramanian et al. (2024) finds that this kind of head specialization also exists in other vision models beyond CLIP. Basile et al. (2024) shows that using only a few selected heads chosen by spectral methods can even beat the full model on zero-shot tasks.
@@ 3 Methodology
In this work, we aim to reduce the activation of redundant attention heads without increasing the number of parameters. A high-level comparison between the vanilla multi-head attention and our proposed MoH is presented in Fig. 1.
@@ 3.1 Multi-Head Attention
We begin by reviewing the multi-head attention mechanism introduced by Vaswani et al. (2017). The multi-head attention mechanism is based on scaled dot-product attention. Specifically, for $T$ tokens $\bm{X}\in\mathbb{R}^{T\times d_{in}}$ of $d_{in}$ dimensions each and $T^{\prime}$ tokens $\bm{X}^{\prime}\in\mathbb{R}^{T^{\prime}\times d_{in}}$ of $d_{in}$ dimensions each, the scaled dot-product attention is computed as follows:
@@
Concatenation Form.    To enhance the representation power, Vaswani et al. (2017) proposes to allow multiple attention heads to operate on different low-dimensional projections of the input tokens. Specifically, the multi-head attention mechanism computes $h$ different low-dimensional projections of $(\bm{Q},\bm{K},\bm{V})$, performs scaled dot-product attention for each head, concatenates the results, and applies a projection to the concatenated output. The concatenation form of the multi-head attention can be formulated as:
@@
where $\bm{W}_{Q}^{i}\in\mathbb{R}^{d_{in}\times d_{k}/h}$, $\bm{W}_{K}^{i}\in\mathbb{R}^{d_{in}\times d_{k}/h}$, and $\bm{W}_{V}^{i}\in\mathbb{R}^{d_{in}\times d_{v}/h}$ represent the $i_{th}$ projection matrices for the query, key, and value, respectively. $\bm{W}_{O}\in\mathbb{R}^{d_{v}\times d_{out}}$ is the final output projection matrix.
@@
Table 1: Comparisons to current state-of-the-art methods on ImageNet-1K classification. Our MoH-ViT models, based on TransNeXt (Shi, 2024), are trained for 300 epochs using a resolution of 224$\times$224. To ensure a fair comparison, we only replace the standard multi-head attention with our Mixture-of-Head attention (MoH), keeping all other training parameters identical to TransNeXt.
@@
3.2 Mixture-of-Head Attention
Recently, the Mixture-of-Experts (MoE) method has emerged as a popular approach for scaling the parameters of large language models (Jiang et al., 2024; Muennighoff et al., 2024). A MoE layer consists of multiple expert networks and a router that activates the Top-K experts. Generally, the number of activated experts $K$ is significantly smaller than the total number of experts to ensure inference efficiency.
Heads as Experts.    Inspired by the great success of MoE, we propose Mixture-of-Head attention (MoH), which treats attention heads as experts. Specifically, MoH consists of $h$ heads $\bm{H}=\{H^{1},H^{2},...,H^{h}\}$ and a router that activates the Top-K attention heads. Formally, given input tokens $\bm{X}$ and $\bm{X}^{\prime}$, the output of MoH is the weighted sum of outputs from the $K$ selected attention heads:
where $g_{i}$ represents the routing score. $g_{i}$ is non-zero only when the $i_{th}$ attention head is activated. This design provides two key advantages: (i) On the one hand, MoH enables each token to select the most relevant attention heads, boosting inference efficiency while maintaining accuracy. (ii) On the other hand, in contrast to the standard summation in multi-head attention, the weighted summation in MoH enhances the flexibility of the attention mechanism and unlocks performance potential.
$\quad g_{i}=\begin{cases}\alpha_{1}\text{Softmax}(\bm{W}_{s}\bm{x}_{t})_{i},&\text{if}\ 1\leq i\leq h_{s},\\ \alpha_{2}\text{Softmax}(\bm{W}_{r}\bm{x}_{t})_{{i-h_{s}}},&\text{if Head $i$ is activated},\\ 0,&\text{otherwise},\end{cases}$ (5)
where $h_{s}$ denotes the number of shared heads. $\bm{W}_{s}\in\mathbb{R}^{h_{s}\times d_{in}}$ and $\bm{W}_{r}\in\mathbb{R}^{(h-h_{s})\times d_{in}}$ represent the projection matrices for the shared and routed heads, respectively. If $(\bm{W}_{r}\bm{x}_{t})_{{i-h_{s}}}\in\text{Top-K}\big{(}\{(\bm{W}_{r}\bm{x}_{t})_{{i-h_{s}}}|h_{s}+1\leq i\leq h\}\big{)}$, then the routed Head $i$ is activated. The coefficients $\alpha_{1}$ and $\alpha_{2}$ balance the contributions of the shared and routed heads, and are defined as:
$[\alpha_{1},\alpha_{2}]=\text{Softmax}(\bm{W}_{h}\bm{x}_{t}),$ (6)
Table 2: Comparisons to current state-of-the-art methods on the benchmarking of class-conditional image generation on ImageNet-1K at 256$\times$256 resolution.$\uparrow$” denotes that higher is better. $\downarrow$” denotes that lower is better. “cfg” denotes the classifier-free diffusion guidance scale. “400K” denotes the training budget is 400K training steps.
@@ 3.2 Mixture-of-Head Attention
DiT-XL/2 7, [^0] K (Peebles & Xie, 2023) 675 100 9.62 6.85 121.50 0.67 0.67
DiT-XL/2 7, [^0] K (cfg=1.25) 675 100 3.22 5.28 201.77 0.76 0.62
MoH-DiT-XL/2 2, [^0] K 676 75 10.95 6.19 106.69 0.67 0.66
MoH-DiT-XL/2 2, [^0] K 676 90 10.67 6.15 107.80 0.67 0.65
MoH-DiT-XL/2 7, [^0] K 676 90 8.56 6.61 129.54 0.68 0.67
MoH-DiT-XL/2 7, [^0] K (cfg=1.25) 676 90 2.94 5.17 207.25 0.77 0.63
$\mathcal{L}_{b}=$ $\sum_{i=h_{s}+1}^{h}P_{i}f_{i},\ P_{i}=\frac{1}{T}\sum_{t=1}^{T}\text{Softmax}(\bm{W}_{r}\bm{x}_{t})_{{i-h_{s}}},$ (7)
$f_{i}=\frac{1}{T}\sum_{t=1}^{T}\mathds{1}(\text{Token $\bm{x}_{t}$ selects Head $i$}),$
$\mathcal{L}=\mathcal{L}_{task}+\beta\mathcal{L}_{b},$ (8)
4 Experiments
4.1 ViT for Image Classification